OpenStudy (anonymous):

Simplify the expression. x^2 - x / 2x - 2x^2 A. –1 B. -1/2 C. 1/2 D. It can’t be simplified.

OpenStudy (anonymous):

@charliealbert

OpenStudy (kkutie7):

Factor out an x from both top and bottom x(x-1)/x(2-2x). (X-1)/(2-2x) Factor out 2 on the bottom (X-1)/2(1-x) and a negative (x-1)/-2(-1+x) cancel x-1 you have -1/2

OpenStudy (anonymous):

\[\frac{ x^{2}-x }{ 2x-x^{2}}\] \[\frac{ x(x-1) }{ -2x(x-1)}\] \[-\frac{ 1 }{ 2 }\]